OpenStudy (anonymous):

Convert 65° into radians

OpenStudy (anonymous):

\[180 \deg = \pi rad => 1 \deg = (\pi \times 1/180) rad => 65 \deg = (\pi \times 65/ 180) rad\]

ganeshie8 (ganeshie8):

\(65 \ degrees \times \frac{\pi\ radians}{180 \ degrees} \)